OpenStudy (anonymous):

A homeowner wants to fence a rectangular garden using 60ft. of fencing. An existing stone wall will be used as one side of the rectangle. Find the dimensions for which the area is a maximum.

OpenStudy (anonymous):

since its a rectangle 2x+y= 60ft, as 1 side is a wall the area of a rectangle is calculated by : xy Isolates y from 2x+y = 60 we have y = 60-2x Subsitute y = 60-2x for xy we have x(60-2x) = 60x-2x^2 Completing the square we (-2x^2+60x) -> -2(x^2-30x) -> -2(x-15)^2+450 Then the maximum point will be (15, 450) Thus x = 15, y = 30 as 2x+y = 60 Maximum area is 450ft squqre

OpenStudy (anonymous):

since its a rectangle 2x+y= 60ft, as 1 side is a wall the area of a rectangle is calculated by : xy Isolates y from 2x+y = 60 we have y = 60-2x Now the Calculus way... Set the Area as a function of x, and substitute what you got for y in the other equation A(x)=x(60-2x) A(x)=60x-2x^2 Now differentiate-- A'(x)=60-4x And set A'(x)=0 0=60-4x 4x=60 x=15 To determine if it's the max, find the second derivative. A"(x)=-4x A"(15)=-4*15 A"15)=-60 Since the second derivative is negative, you know the function is concave down at that point, making 15 a maximum. Now use the original information to find y.